Bob Agrajag^{Hitchhiker’s Guide: B Ark}_{Character Sheet}

Arthur Dent … I hate that guy! You

You really hate Arthur Dent. You’re not sure why you know that, or even who Arthur Dent is, but you’re sure that you hate him. You see, right now you’re Robert Agrajag, cleanliness cechnician^{1} for the Golgafrinchan Ark Fleet, Ship B, but you remember being many other things. A London executive spending an evening in a nightclub. A rabbit in a peaceful forest. Even a bowl of petunias mysteriously called into existance miles up into the air. You have been reincarnated in many different forms, and in many different times. Unlike most, though, you can remember your previous incarnations.

The reason for this is that all of these memories have one thing in common. All of them include being killed, and each time it was Arthur Dent’s fault. Oh, he didn’t always do it on purpose, though he did knock your rabbit-self over the head for dinner. Sometimes, like at… Stavromula Beta was it? The memories aren’t too clear… Sometimes he was just ducking a bullet and it happened to hit you. In your petunia-self you just saw him looking out the window of a passing ship, you can’t be sure if he had anything to do with your rapid demise. Despite the immediate evidence, though, the pattern is clear. Arthur Dent exists to make your life miserable, and thus you must destroy him.

Of course, you never know when Arthur Dent may appear, so you just have to go on with your life until then. You rather like being a cleanliness technician. It’s quiet work and you’re left much to yourself, especially on the Ark where you’re the only one around. Keeping all of the freezer bays clean isn’t really that tough with nobody tracking dirt on the �oor after all.

Bluesheets

• Society of Cleanliness Engineers Greensheets

• Cleaning Schedules Abilities

• Sample Powers Items

• A Mop.

A nameplate on the handle reads �Bob Agrajag�. This is your mop. Your daddy gave it to you when you were eight years old.

• A Golgafrinchan ID Card for �Bob Agrajag�. Opens type C locks.

• A bucket.

Holds 2 gallons of water.

Memory Packets

• Badge number 42

• Open at Game Start Stats • Combat Ability (ϕ): 5 • Alpha (α): 2

1Read janitor.